OpenStudy (anonymous):

A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?

OpenStudy (anonymous):

first separate x from y -7+3y=-21.5/x next, try and move all the numbers from the y side to the x side 3y=(-21.5/x)+7 y=([-21.5/x]+7)/3 now solve